word2:

Determine whether the vectors u and v are parallel, orthogonal, or neither. u = <9, 3>, v = <36, 12> Parallel Orthogonal Neither can someone check my work ? i think the answer is neither because it doesn't equal to zero and they aren't equal when you multiply it. is this correct?

word2:

@theDeviliscoming

theDeviliscoming:

I'm so sorry but I am horrible at math maybe @Hero can help

Hero:

I'm about to go for a run. I probably should not have come here. It is not difficult to find videos or websites that show how to solve problems like these.

word2:

@Nnesha Can you quickly check this for me?

Nnesha:

what isn't equal to zero ??

word2:

for it to be orthogonal u ▪ v = 0

Nnesha:

yes right

word2:

in this problem u ▪ v = 360 so i think its neither because u = 324 , v = 36

Nnesha:

did you find the angle between them ?

Nnesha:

ok

word2:

no , i didn't think i needed to do that because parallel means they're the same correct?

Nnesha:

If the angle between them is zero or 180 then it will be parallel

word2:

okay so then its the LAST one?

Nnesha:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @word2 no , i didn't think i needed to do that because parallel means they're the same correct? \(\color{#0cbb34}{\text{End of Quote}}\) yes but what exactly should be the same ? \[\large\rm u \cdot v \color{reD}{= }\sqrt{u} \sqrt{v}\] if the above equation is true then it will be parallel

word2:

108?

Nnesha:

hmm what is the magnitude of u and v ?

word2:

\[\sqrt{324}\sqrt{36} = 108\]

Nnesha:

the x-component of vector v is 36 so the magnitude should be bigger than that

word2:

\[\sqrt{9^2+ 36^2}= 37.01\]

Nnesha:

\

Nnesha:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Nnesha \ \(\color{#0cbb34}{\text{End of Quote}}\) \[\large\rm u=<\color{Red}{9},\color{blue}{3}> ~~~\rightarrow \left| u \right|= \sqrt{\color{reD}{9^2}+\color{blue}{3^2} }\] https://prnt.sc/g422dq

Nnesha:

the component of vector v are 36 and 12 ( not 36 and 9 ) :=))

word2:

i did that , 9 * 36 = 324 , 3 *12 = 36

word2:

thanks eveyone , i;m going to close this . i go it

Nnesha:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @word2 in this problem u ▪ v = 360 so i think its neither because u = 324 , v = 36 \(\color{#0cbb34}{\text{End of Quote}}\) dot product is 360 but we cannot separate them like u =324 , v=36 \

Nnesha:

\[ \large\rm u \cdot v = 324+36\] **

Nnesha:

\[\large\rm u=<\color{Red}{9},\color{blue}{3}> ~~~\rightarrow \left| u \right|= \sqrt{\color{reD}{9^2}+\color{blue}{3^2} }\] \[\large\rm v=<\color{Red}{36},\color{blue}{12}> ~~~\rightarrow \left| v \right|= \sqrt{\color{reD}{(36)^2}+\color{blue}{(12)^2} }\] \[\large\rm u \cdot v = \left| u \right|\left| v \right|\]

Nnesha:

are you talking to me ?

Nnesha:

hmm

Nnesha:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @sillybilly123 \(\mathbf u = <9,3>\) \(\mathbf v = <36,12> = 4 <9,3>\) scalar linearity :) \(\color{#0cbb34}{\text{End of Quote}}\) agree w/ ya

Nnesha:

cool o^_^o

sillybilly123:

o^_^o